How many electrons, protons, and neutrons in ^40 CA ^+2

Please show the answer step by step or show the reason for the number you choose! Thanks
Answer #1

Ans: The atomic number of Calcium ( Ca ) is 20

Atomic Number = Number of Electrons = Number of Protons = 20.

. Now,  electronic configuration of Ca and Ca2+ is:

Ca : 1s2 2s2 2p6 3s2 3p6 4s2.

Ca2+ : 1s2 2s2 2p6 3s2 3p6

Here, Ca2+ ion has 2 electron less.

So, Number of electrons , e = 20 - 2 = 18

Number of protons , p = Atomic Number = 20 .

Number of Neutrons , n = Mass Number - Atomic Number = 40 - 20 = 20.